FOOTPRINT92: Logbook Entries

FOOTPRINT92: PSYC Messages: 8 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
14 Thu 04-Jun-1992PRESSURE SENSOR MOVED: nonesrs
53 Mon 08-Jun-1992PRESSURE SENSOR BAD: nonesrs
85 Thu 11-Jun-1992PSYC TOWER COMING DOWN: nonesrs
86 Thu 11-Jun-1992CHANGING THE PSYCHROMETERS noneacd
90 Thu 11-Jun-1992PSYC @ 3M moved to 10M: nonesrs
91 Thu 11-Jun-1992INSTALL PSYC #14 @ 3M: nonesrs
104 Sun 14-Jun-19923m Psychrometer nonetwh
113 Tue 16-Jun-19923m Psychrometer humidity mismatched to nonetwh

14: PSYC , Site none, Thu 04-Jun-1992 18:57:18 GMT, PRESSURE SENSOR MOVED:
	Due to software conditions the pressure 
sensor was moved back to the 250m tower, cosmos. 
The conditional sampler system will be setup at 
the 50m tower, ragwort. 
53: PSYC , Site none, Mon 08-Jun-1992 00:53:13 GMT, PRESSURE SENSOR BAD:
	The pressure sensor has not been 
working too good. It has a tendency to go 
down everytime you do anything with cosmos. 
The sensor is down at this time. 
85: PSYC , Site none, Thu 11-Jun-1992 15:11:44 GMT, PSYC TOWER COMING DOWN:
	The psyc tower will be down for awhile 
to replace the psycs with calibrated humidity 
86: PSYC , Site none, Thu 11-Jun-1992 16:50:59 GMT, CHANGING THE PSYCHROMETERS
Changing the psychrometers 
At approx 08:00 the psych tower was lowered and the 
psychrometers were removed. They were replaced by new ones. 
                          Old       New 
	        10.0m     #67       #14 
		 7.5m     #21       #13 
		 5.0m     #36       #59 
		 3.0m     #23       #28 
		 1.5m     #35       #65 
90: PSYC , Site none, Thu 11-Jun-1992 21:19:00 GMT, PSYC @ 3M moved to 10M:
PSYC @ 3M moved to 10M: 
	The psyc at 10m, S/N 14, was giving 
low humidity readings. It was decided to move 
the 3m unit, S/N 28 up to the 10m height. 
S/N 14 will be looked at for incorrect coefficients. 
This change took place about 15 minute ago. 
91: PSYC , Site none, Thu 11-Jun-1992 22:08:38 GMT, INSTALL PSYC #14 @ 3M:
	Psyc #14 was installed at the 3m 
mark. I could not get the unit to respond to 
commands to check the coefficients. Suggest replacing 
this unit with the spare one if TOm has the 
humitter for it. 
104: PSYC , Site none, Sun 14-Jun-1992 00:26:54 GMT, 3m Psychrometer
3m Psychrometer 
Replaced current 3m psychrometer (14) with 
spare psychrometer (34) after installing 
humitter that Cathy sent from cal lab. 
The relative humidity at 3m may still be 
a little low  (or 1.5m may be high!). 
113: PSYC , Site none, Tue 16-Jun-1992 16:05:45 GMT, 3m Psychrometer humidity mismatched to
3m Psychrometer humidity mismatched to others. 
The specific humidity data at 3m does not match that 
at the other levels.  It is offset and in some 
cases appears not as responsive.  I checked 
the fan and it is running okay. 
Note added at 01:25 GMT, June 17:  However, the 
relative humidity profile looks believable.  I 
think that the specific humidity profile is  
questionable because we are near the measurement 
uncertainty for that variable.  The specific 
humidity differences are less than 0.1 gm/kg.